A report gave the following information on births in a certain country during a particular year.

Type of Birth Number of Births
Single birth 3,838,214
Twins 125,336
Triplets 4,233
Quadruplets 236
Quintuplets or higher 47

(a)

Use this information to estimate the probability that a randomly selected pregnant woman who gave birth in this year delivered twins. (Round your answer to three decimal places.)

(b)

Use this information to estimate the probability that a randomly selected pregnant woman who gave birth in this year delivered quadruplets. (Round your answer to six decimal places.)

(c)

Use this information to estimate the probability that a randomly selected pregnant woman who gave birth in this year gave birth to more than a single child. (Round your answer to three decimal places.)
Answer #1

There are total number of births= 3968066


a)

P(twins) = 125336/3968066

= 0.032


b)

P(quadruplets) = 236/3968066
= 0.000059


c)

P(more tan a single child) = (125336 + 4233 + 236 + 47)/3968066
= 0.033